Vector A has a magnitude of 146 units and points 36.0° north of west. Vector B points 63.0° east of north. Vector points 16.0° west of south. These three vectors add to give a resultant vector that is zero. Using components, find the magnitudes of the following vectors e) vector B x units (b) vector C x units Enter a number
Answer #1


\a_X=-146;cos36\\ A_y=146;sin36\\ B_x=B;sin63\\ B_y=B;cos63\\ C_x=-C;sin16\\ C_y=-C;cos16\\ given;resultant;vector=0\\ R_x=0\\ A_x + B_x + C_x=0\\ -146;cos36+Bsin63-Csin16=0\ Csin16=-146;cos36+B;sin63...(1)\\\ R_y=0\\ A_y + B_y + C_y=0\\ 146;sin36+Bcos63-Ccos16=0\ Ccos16=146;sin36+B;cos63.....(2)\\\from;(1);and;(2)\ tan16=rac{-146;cos36+B;sin63}{146;sin36+B;cos63}\\ B=188;\\ using;B;in;1\ ax--146 cos36 Ay 146 sin36 By-B cos63 CT-C sin16 C,--C cos16 given resultant vector0 146 cos36+ Bsin63 - Csin 16 0 Csin16-146

using B in (1) C 179.2